Suppose that we are testing the null hypothesis that μ = 68 versus the alternative that μ < 68. We take a sample of size 36 and find a sample average of = 63.4 and a sample standard deviation of s = 12.6. Determine the value of the test statistic for the hypothesis test of one population mean. t = -2.19

t = -0.37

t = 0.37

t = 2.19
Answer #1

Test statistic = t

= (\bar x - \mu ) / s / \sqrt n

= (63.4-68 ) / 12.6 / \sqrt 36

t= -2.19

Test statistic = t= - 2.19